OpenStudy (anonymous):

Can someone please help me? I have no idea how I need to set this up. The average price of a movie ticket in the United States from 1980 to 2000 can be modeled by the function f(x) = 0.10x + 2.75 where x is the number of years since 1980. A. Graph the function and identify its domain and range. B. Find the vale of X so that f(x) = 4.55 and explain what the solution means in this situation.

OpenStudy (trojanpoem):

A) I think the domain must remain positive in this case as f(x) the average may not be negative. f(x) in (inf, 0] (range) 0.10x + 2.75 >= 0 0.10x >= -2.75 x >= -27.5 x in (inf, -27.5] (domain) (Not sure of first part indeed) 0.10 x + 2.75 = 4.55 x = 18 ( x is the number of years since 1980 ) year -> 1998 the average ticket price was 4.55